# Building Trustworthy AI: Transparent AI Systems via Large Language Models, Ontologies, and Logical Reasoning (TranspNet)

### v1

@@ Abstract
Growing concerns over the lack of transparency in AI, particularly in high-stakes fields like healthcare and finance, drive the need for explainable and trustworthy systems. While Large Language Models (LLMs) perform exceptionally well in generating accurate outputs, their "black box" nature poses significant challenges to transparency and trust. To address this, the paper proposes the TranspNet pipeline, which integrates symbolic AI with LLMs. By leveraging domain expert knowledge, retrieval-augmented generation (RAG), and formal reasoning frameworks like Answer Set Programming (ASP), TranspNet enhances LLM outputs with structured reasoning and verification. This approach ensures that AI systems deliver not only accurate but also explainable and trustworthy results, meeting regulatory demands for transparency and accountability. TranspNet provides a comprehensive solution for developing AI systems that are reliable and interpretable, making it suitable for real-world applications where trust is critical.
@@ 1 Introduction
By combining the power of LLMs with formal reasoning, retrieval-augmented generation, multimodal data processing, and robust documentation practices, our pipeline addresses the key challenges associated with explainability and trustworthiness in AI systems. The integration of formal logic through ASP further distinguishes our approach by providing a mechanism for verifying the logical soundness of the LLM’s outputs and addressing uncertainty in a structured and interpretable manner. Ultimately, our pipeline offers a comprehensive solution for developing LLM-based AI systems that are not only accurate but also explainable and trustworthy, meeting the needs of both academic research and industry applications.
@@ 2 State-of-the-Art in LLM Reasoning
The proposed pipeline enhances the state-of-the-art by addressing some of the core challenges faced by LLMs, particularly in ensuring explainability and trustworthiness in reasoning tasks. Unlike traditional LLM systems, the pipeline integrates domain expert knowledge, retrieval-augmented generation, and formal reasoning frameworks, enabling the system to verify and refine its outputs through external data and structured reasoning. This multi-layered approach mitigates some of the inherent issues in LLMs, such as hallucination and over-reliance on probabilistic reasoning. The inclusion of ASP [13] in the pipeline ensures that the reasoning process is grounded in logical consistency, providing verifiable and interpretable outputs. By incorporating multimodal data processing, the pipeline allows the system to handle diverse types of input, further improving its reliability and decision-making capabilities across different domains.
@@ 3 Methodology
Figure 1: Overall knowledge-driven LLM verification pipeline. The pipeline integrates domain expert knowledge, prompt engineering, retrieval-augmented generation, answer set programming, and deep learning models to ensure accurate, explainable, and trustworthy outputs from LLMs.
@@ 3.1 Overall Knowledge-Driven LLM Verification Pipeline
This pipeline integrates domain expert knowledge and structured data to enhance the reliability and LLM outputs. The process starts with expert knowledge and vocabulary analysis, followed by prompt engineering and RAG to leverage large language models for generating answers. Ontology matching ensures semantic consistency, with structured ontologies being used for mapping concepts. A critical component of the pipeline is the consciousness layer for LLMs, which incorporates ASP for applying logical reasoning. This layer enhances the verification, explanation, and trustworthiness of the generated answers, leading to a robust system aimed at providing accurate and explainable outputs [2, 9].
@@
The pipeline begins with the input from domain experts and relevant structured data. These inputs ensure the accuracy and relevance of the information. The data may include technical documents, research papers, and domain-specific databases.
@@
The identification of vocabulary, attributes, and relationships is a crucial step in the pipeline. By defining the key terms, their attributes, and relationships, this foundational phase ensures that the LLM understands the specific context and nuances of the subject matter. This step enables accurate and meaningful interactions in subsequent stages, effectively mapping domain knowledge into a structured format to improve the comprehension and performance of the LLM.
@@
Once the input query is refined through prompt engineering, the LLM generates structured responses in a subject-predicate-object format. The LLM’s extensive training on diverse datasets enables it to produce contextually relevant responses that maintain consistency with the controlled vocabulary defined in earlier steps. These generated triples are then mapped to the ontology, facilitating further reasoning and analysis in the pipeline.
@@
Concept matching compares elements in the generated triples with those in the ontology to ensure alignment. Multiple matching techniques are used, including: - Name-Based Matching: String matching algorithms like Levenshtein distance and Jaccard similarity to compare element names [11]. - Structure-Based Matching: Examination of class hierarchies and property relationships in the schema [21]. - Instance-Based Matching: Comparison of actual data instances for value similarity [8]. - Linguistic Matching: Use of natural language processing techniques, such as synonym databases and word embeddings, to find semantically similar matches [31].
@@
The role of the Consciousness Layer is to reason on the concepts extracted from (2.a) and the structured response from LLM (1.e). Therefore, the required piece of information into the domain specific ontology will be use as a knowledge-base for the ASP solver. The Consciousness Layer includes: - ASP Knowledge Base: Stores factual and procedural knowledge from mapped concepts and domain knowledge. - ASP Rules: Logical rules and constraints for reasoning over the knowledge base. - ASP Solver: The ASP solver applies logical inference to refine LLM-generated answers, ensuring consistency, accuracy, and robustness. This solver verifies that the generated triples are logically sound and contextually relevant [12].
@@
Extracted features refer to the characteristics obtained after applying deep learning models for tasks such as classification, regression, and clustering. These models process data from various multimodal sources, including sensors, to extract meaningful and relevant features. This approach is particularly advantageous for our use cases, especially in the Chemical Processes context, where precise feature extraction from multimodal sensor data is crucial for accurate analysis and prediction. To further enhance the interpretability and utility of these extracted features, it is important to map them to higher-level concepts. Techniques such as the bottleneck model, inspired by the DeVil model, can be employed to identify and distill key features that contribute most significantly to the model’s performance. By identifying these critical features and relating them to specific concepts, we can achieve a more intuitive understanding of the model’s decision-making process.
@@ 5 Conclusion
In this work, we presented a robust and comprehensive pipeline designed to enhance the explainability, accuracy, and trustworthiness of LLMs. By integrating domain expert knowledge, prompt engineering, RAG, and ASP, the proposed pipeline addresses key challenges in ensuring that LLM-generated outputs are logically consistent, contextually relevant, and semantically aligned with domain-specific knowledge.
The pipeline is structured to ensure that each stage of the process, from vocabulary identification to ontology mapping and logical reasoning, contributes to producing reliable outputs. The Consciousness Layer, with its ASP-based reasoning, represents a novel advancement in validating LLM outputs, offering a higher level of trust and transparency by providing verifiable logical consistency.
The practical application of the pipeline in diverse domains, such as healthcare and battery design, demonstrates its versatility and effectiveness in generating evidence-based, explainable, and trustworthy recommendations. In healthcare, the pipeline ensures that clinical decision support systems produce recommendations aligned with medical guidelines, ensuring patient safety and improving the quality of care. In battery design, the pipeline helps engineers make informed material selections, optimizing performance and longevity in energy storage systems.
Overall, the proposed pipeline is a significant step forward in addressing the limitations of traditional LLM outputs. It offers a comprehensive framework for generating not only accurate and context-aware answers but also ensuring that these answers are explainable, reliable, and suitable for real-world applications. As LLMs continue to evolve and are increasingly adopted across various domains, the integration of structured reasoning and expert knowledge, as demonstrated in this pipeline, will be crucial in maintaining trust and ensuring the success of AI-driven systems.

### v2

# Building Trustworthy AI: Transparent AI Systems via Language Models, Ontologies, and Logical Reasoning (TranspNet)

@@ Abstract
Growing concerns over the lack of transparency in AI, particularly in high-stakes fields like healthcare and finance, drive the need for explainable and trustworthy systems. While Large Language Models (LLMs) perform exceptionally well in generating accurate outputs, their "black box" nature poses significant challenges to transparency and trust. To address this, the paper proposes the TranspNet pipeline, which integrates symbolic AI with LLMs. By leveraging domain expert knowledge, retrieval-augmented generation (RAG), and formal reasoning frameworks like Answer Set Programming (ASP), TranspNet enhances LLM outputs with structured reasoning and verification.This approach strives to help AI systems deliver results that are as accurate, explainable, and trustworthy as possible, aligning with regulatory expectations for transparency and accountability. TranspNet provides a solution for developing AI systems that are reliable and interpretable, making it suitable for real-world applications where trust is critical.
Keywords:
Trustworthy AI Explainable AI Large Language Models (LLMs) Symbolic AI Ontologies Answer Set Programming (ASP) Retrieval-Augmented Generation (RAG) AI Transparency AI Accountability Neural-Symbolic Integration.
@@ 1 Introduction
By combining the power of LLMs with formal reasoning, retrieval-augmented generation, multimodal data processing, and robust documentation practices, our pipeline addresses the key challenges associated with explainability and trustworthiness in AI systems. The integration of formal logic through ASP further distinguishes our approach by providing a mechanism for verifying the logical soundness of the LLM’s outputs and addressing uncertainty in a structured and interpretable manner.
@@ 2 State-of-the-Art in LLM Reasoning
The proposed pipeline integrates domain expert knowledge, retrieval-augmented generation, and formal reasoning frameworks, enabling the system to verify and refine its outputs through external data and structured reasoning. This multi-layered approach mitigates some of the inherent issues in LLMs, such as hallucination and over-reliance on probabilistic reasoning. The inclusion of ASP [13] in the pipeline ensures that the reasoning process is grounded in logical consistency, providing verifiable and interpretable outputs. By incorporating multimodal data processing, the pipeline allows the system to handle diverse types of input, further improving its reliability and decision-making capabilities across different domains.
@@ 3 Methodology
Figure 1: Overall knowledge-driven LLM verification pipeline. The pipeline integrates domain expert knowledge, prompt engineering, retrieval-augmented generation, answer set programming, and deep learning models to support the generation of accurate, explainable, and trustworthy outputs from LLMs.
@@ 3.1 Overall Knowledge-Driven LLM Verification Pipeline
The proposed pipeline integrates domain expert knowledge and structured data to enhance the reliability and LLM outputs. The process starts with expert knowledge and vocabulary analysis, followed by prompt engineering and RAG to leverage large language models for generating answers. Ontology matching ensures semantic consistency, with structured ontologies being used for mapping concepts. A critical component of the pipeline is the consciousness layer for LLMs, which incorporates ASP for applying logical reasoning. This layer enhances the verification, explanation, and trustworthiness of the generated answers, leading to a robust system aimed at providing accurate and explainable outputs [2, 9].
@@
The pipeline begins with the input from domain experts and relevant structured data. The data may include technical documents, research papers, and domain-specific databases.
@@
The identification of vocabulary, attributes, and relationships is a foundational step in the pipeline, ensuring semantic consistency and alignment with domain-specific knowledge. This process involves defining key terms (vocabulary), their properties (attributes), and the logical or causal connections between them (relationships). For example, in healthcare, vocabulary might include terms like "symptom," "diagnosis," and "treatment," with attributes such as severity or duration, and relationships such as "symptom X is linked to diagnosis Y." This step can be performed manually by domain experts, who ensure precision and relevance, or automated using tools like natural language processing and machine learning to extract terms and relationships from structured datasets or research papers. Often, a hybrid approach is used, where automated methods handle initial extraction, and experts refine or validate the results.
@@
Large Language Models (LLMs) are employed for additional tasks such as refining responses, summarization, and generating structured outputs.
In this pipeline, Retrieval-Augmented Generation (RAG) first retrieves relevant context from external sources. This context is then provided to the LLM to guide its response generation, ensuring outputs are grounded in reliable and contextually relevant information.
After the input query is refined through prompt engineering, the LLM generates structured responses in a subject-predicate-object format. Leveraging its extensive training on diverse datasets, the LLM produces contextually accurate outputs aligned with the controlled vocabulary defined in earlier steps. These generated triples are subsequently mapped to the ontology, enabling advanced reasoning and analysis in downstream processes.
@@
Concept matching compares elements in the generated triples with those in the ontology to ensure alignment. Multiple matching techniques are used, including:
  • Name-Based Matching: String matching algorithms like Levenshtein distance and Jaccard similarity to compare element names [11].
  • Structure-Based Matching: Examination of class hierarchies and property relationships in the schema [21].
  • Instance-Based Matching: Comparison of actual data instances for value similarity [8].
  • Linguistic Matching: Use of natural language processing techniques, such as synonym databases and word embeddings, to find semantically similar matches [31].
@@
The role of the Consciousness Layer is to reason on the concepts extracted from (2.a) and the structured response from LLM (1.e). Therefore, the required piece of information into the domain specific ontology will be use as a knowledge-base for the ASP solver. The Consciousness Layer includes:
  • ASP Knowledge Base: Stores factual and procedural knowledge from mapped concepts and domain knowledge.
  • ASP Rules: Logical rules and constraints for reasoning over the knowledge base.
  • ASP Solver: The ASP solver applies logical inference to refine LLM-generated answers, ensuring consistency, accuracy, and robustness. This solver verifies that the generated triples are logically sound and contextually relevant [12].
@@
Extracted features refer to the characteristics obtained after applying deep learning models for tasks such as classification, regression, or clustering. These models process data from various multimodal sources, including sensors, to extract meaningful and relevant features. This approach is particularly advantageous for our use cases, for example in the Chemical Processes context, where precise feature extraction from multimodal sensor data is crucial for accurate analysis and prediction. To further enhance the interpretability and utility of these extracted features, it is important to map them to higher-level concepts. Techniques such as the bottleneck model, inspired by the DeVil model, can be employed to identify and distill key features that contribute most significantly to the model’s performance. By identifying these critical features and relating them to specific concepts, we can achieve a more intuitive understanding of the model’s decision-making process.
@@ 4.2 Use Case 2: Battery Design - Material Selection for Energy Storage
5 Limitations
One limitation of the proposed pipeline is that its emphasis on structured reasoning and logical consistency might reduce the flexibility and adaptability of LLMs in certain contexts. For example, in applications like healthcare or education, the pipeline excels by ensuring outputs are accurate, reliable, and aligned with established guidelines. This is especially important when making decisions that directly impact patient safety or the clarity of learning materials.
However, this structured approach might not be as effective in more creative or exploratory applications, such as marketing, content creation, or brainstorming new product ideas. In these scenarios, flexibility, innovation, and the ability to generate unconventional or imaginative responses are often more valuable than strict logical accuracy. By limiting the LLM’s predictive power to fit within predefined rules and structures, the pipeline may constrain its potential to explore diverse or unexpected outcomes.
@@ 6 Conclusion
In this work, we presented a pipeline designed to enhance the explainability, accuracy, and trustworthiness of LLMs. By integrating domain expert knowledge, prompt engineering, RAG, and ASP, the proposed pipeline addresses key challenges in ensuring that LLM-generated outputs are logically consistent, contextually relevant, and semantically aligned with domain-specific knowledge.
The proposed pipeline represents an advancement in addressing the inherent limitations of traditional LLMs. Its ASP-based Consciousness Layer sets a benchmark for ensuring logical consistency, enhancing trust, and fostering greater transparency in AI outputs. As LLMs continue to evolve and find applications in increasingly critical domains, approaches that combine structured reasoning and expert knowledge, as exemplified by this pipeline, will be essential for reliability, transparency, and success in real-world implementations.
{credits}
6.0.1 Acknowledgements
Funding is acknowledged from the European Union’s Horizon Europe research and innovation programme under grant agreements no. 101137725 (BatCAT) and 101138510 (DigiPass CSA).
[…]
The authors have submitted a grant proposal in which the presented architecture plays a central role.